OpenStudy (anonymous):

10m^2+2m-3=0 use the quadratic formula

OpenStudy (anonymous):

i wanna make sure i did it correct

OpenStudy (anonymous):

plug and chug. \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

ok we get a = 10 b = 2 c = -3

OpenStudy (anonymous):

best to write them down and then you know what you are substituting

OpenStudy (anonymous):

less likely to screw up

OpenStudy (anonymous):

What is the answer you got? We can tell you if its right or not, This is the 5th one of these ive seen from you :P maybe we can teach you instead of give answers. Again, it is in standard form of \[ax^2 + bx + c = 0\] now, just plug and chug into the quadratic equation :)

OpenStudy (anonymous):

\[\frac{-2\pm\sqrt{2^2-4\times 10\times -3}}{2\times 10}\]

OpenStudy (anonymous):

ok so -2+- square root 2^2-4(2)(-3)/2(10)

OpenStudy (anonymous):

i get \[\frac{-1\pm\sqrt{31}}{10}\]

OpenStudy (anonymous):

oh wait i am sorry you are right

OpenStudy (anonymous):

you are correct.

OpenStudy (anonymous):

\[-2\pm \sqrt{28}\]/20

OpenStudy (anonymous):

get \[\frac{-2\pm\sqrt{124}}{20}\]

OpenStudy (anonymous):

think you messed up with \[b^2-4ac\] again

OpenStudy (anonymous):

\[2^2-4\times 10\times -3\] two minus signs so it is plus

OpenStudy (anonymous):

\[4+120\]

OpenStudy (anonymous):

124

OpenStudy (anonymous):

-2 +- square root 124/20